OpenStudy (solomonzelman):

What does it mean to use a unit circle? Where is that applied? ( Explain, no links! )

OpenStudy (anonymous):

A unit circle is a circle with a radius of 1. 

OpenStudy (solomonzelman):

OK! and where do you apply this?

OpenStudy (anonymous):

It's elegant for a lot of applications. Like, we know that \(x^2 + y^2 = 1\) because of Pythagoras. So since we also know that any coordinate on the circle is \((\cos\theta,\sin\theta)\) we can prove the useful identity \(cos^2\theta + sin^2\theta = 1\). 

OpenStudy (solomonzelman):

Can you tell me more about this please?

OpenStudy (anonymous):

Well, sure. Like what would you like to know?

OpenStudy (anonymous):

I mean, the really simple and short answer is "the unit circle is good for proving a lot of really useful math stuff."

OpenStudy (solomonzelman):

An example of how this is useful to real life

OpenStudy (solomonzelman):

U know what mean? Like how is this math practical and necessary.

OpenStudy (anonymous):

It's useful in physics. I don't actually know enough physics to be able to really go into a lot of detail, but physics problems are often a lot more closer to "real life" than math problems... :D

OpenStudy (solomonzelman):

Can you show me how people would apply this knowledge to their real life?

OpenStudy (anonymous):

It depends on what people. A lot of people can make it through their life just fine only knowing how to add and subtract and multiply and divide. A lot of higher math just isn't relevant. That said, for the professionals in a field, it definitely is. For example, if you're just playing a video game, you don't have to care about this, but circles are a really quick and easy way of doing 2d collision detection (i.e., determining if one object bumps into another) so the people who made the game certainly know and care about this math.

OpenStudy (jdoe0001):

SolomonZelman do you have a Unit Circle yourself?

OpenStudy (solomonzelman):

OK, see!